how do i do an equation like this: 9y+1-y=49
Answer by Alan3354(69443)

how do i do an equation like this: 9y+1-y=49
----------
Define "do"
9y+1-y=49
You can solve for y
8y+1 = 49
Subtract 1
8y = 48
Now what?